# CDKFormer: Contextual Deviation Knowledge-Based Transformer for Long-Tail Trajectory Prediction

### v1

@@ III-C Long-Tail Characteristics
Utilizing Argoverse 2 motion forecasting dataset [6], we investigate the long-tail characteristics of the target vehicle trajectories from two perspectives: individual motion and group interaction. Our analysis focuses on comparing the top 10% of tailed samples with the top 10% of head samples, based on the defined tail score. The following metrics are proposed to quantitatively measure the deviations in individual motion and group interactions between tailed and normal scenarios.
@@
These metrics, which represent the complexity of traffic scenarios surrounding the target vehicle, are independent of vehicle absolute positions. Consequently, they can be utilized to effectively assess the deviation of the current traffic state in a coordinate-agnostic way.
A descriptive table on these metrics and the results of significance comparisons from ANOVA tests are summarized in Table I. The distributions of individual and group deviation measures are shown in Fig. 3 and Fig. 4, respectively. As a result, tail samples exhibit significantly higher mean changes in velocity and greater variability in both velocity and heading. For example, among the top 10% of tail samples, 26.74% involve clear turning maneuvers, while the remaining 73.26% proceed straight. In contrast, 88.28% of the top 10% head samples follow straight trajectories. Additionally, tail samples show more dynamic agent interactions with higher relative speed and greater group heading variability. These distinctions underscore the complexity and unpredictability inherent in long-tailed vehicle trajectories, highlighting the necessity of incorporating these features to fully understand and model long-tail traffic scenarios.
TABLE I: Long-tail characteristics of trajectory samples in Argoverse 2 motion forecasting dataset.
@@
Individual $\Delta V_{\rm ind}$ 1.37$\pm$3.59 -0.14$\pm$0.75 $<$.001
$\Delta H_{\rm ind}$ 0.12$\pm$77.07 -0.12$\pm$18.06 .69
$\sigma(V_{\rm ind})$ 1.10$\pm$0.85 0.05$\pm$0.26 $<$.001
$\sigma(H_{\rm ind})$ 13.09$\pm$33.49 0.59$\pm$8.24 $<$.001
Group $\Delta V_{\rm grp}$ 5.59$\pm$3.46 0.79$\pm$1.20 $<$.001
$\sigma(H_{\rm grp})$ 76.58$\pm$42.91 42.00$\pm$47.93 $<$.001
@@
IV Methodology
IV-A Prediction Framework Overview
The proposed CDKFormer framework is illustrated in Fig. 5. CDKFormer is constructed in a encoder-decoder way. In the scene encoder, we first encode the HD map vectors and agent motion information and fuse them in a scene context fusion module. The deviation feature is learned in parallel with a deviation fusion module. Then, the learned context feature and deviation feature are jointly decoded in a query-based paradigm. A mode query and dual future queries are initiated to interactively and progressively extract the context and deviation feature through multistream decoder blocks. Then a scene query is obtained by a learnable gating mechanism. We additionally refine this scene query and generate the final multimodal trajectory predictions.
Figure 5: Overview of the proposed CDKFormer architecture. The model first encode the agent motion and scene contextual information with self-attention-based encoders. The deviation and motion features of the target vehicle are jointly fused in a deviation fusion module. The scene context and deviation information are subsequently decoded by a mode query and dual future queries, including a regular future query and a tail future query, within multistream decoder blocks. Then, a scene query is obtained by combining the mode query and weighted combined future query. This scene query is further refined and used for multimodal trajectories generation. ×$\times N$ denotes $N$ stacked layers. (R)Future and (T)Future denote regular future query and tail future query, respectively.
IV-B Scene Context Encoding
IV-B1 Agent Motion Encoding
The agent motion input of the module is a 6-dimensional vector $\bm{X}\in\mathbb{R}^{N_{a}\times T_{o}\times d_{a}}$, where $d_{a}=6$ includes the 2D historical position, positional displacement vector, positional displacement magnitude, and absolute velocity. Fourier embedding [46] is first applied to these motion inputs, followed by a multilayer perceptron (MLP) to project them into a high-dimensional space, producing the group motion encoding $\bm{X}_{a}\in\mathbb{R}^{N_{a}\times T_{o}\times d}$. Then, we apply a standard Transformer encoder on $\bm{X}_{a}$, which is composed of a multihead attention and an MLP, with layer normalization and residual connections added. This enables feature aggregation in the temporal dimension. The multihead attention is calculated as follows.
@@ IV-B1 Agent Motion Encoding
where $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are weight matrix, $\bm{X}$ is the input, $d_{k}$ is the hidden state dimension. The encoder layer is repeated for $N$ times to obtain the motion feature $\bm{\tilde{X}}_{a}\in\mathbb{R}^{N_{a}\times T_{o}\times d}$, which represents the motion dynamic of both the target vehicle and its surrounding agents. Then we extract the target vehicle’s motion feature $\bm{X}_{\rm tgt}\in\mathbb{R}^{T_{o}\times d}$ from $\bm{\tilde{X}}_{a}$, which will be used for deviation feature fusion in Section IV-C. The final agent motion encoding is obtained by extracting the last timestep of $\bm{\tilde{X}}_{a}$, denoted as $\bm{C}_{a}\in\mathbb{R}^{N_{a}\times d}$.
IV-B2 Map Encoding
The input HD map feature $\bm{I}\in\mathbb{R}^{N_{m}\times l_{m}\times d_{m}}$ comprises positions and displacement vectors of centerlines surrounding the target vehicle, organized into $N_{m}$ polylines, each containing $l_{m}$ points with $d_{m}=4$. We encode these polylines with a PointNet-like encoder [47], as also employed in [20, 48].
@@ IV-B2 Map Encoding
IV-B3 Scene Context Fusion
Scene context fusion is conducted to fully integrate traffic agent motion and map information, enabling a semantic understanding of dynamic traffic environments. We first concatenate $\bm{C}_{a}\in\mathbb{R}^{N_{a}\times d}$ and $\bm{C}_{m}\in\mathbb{R}^{N_{m}\times d}$ to a scene encoding $\bm{C}_{\rm sce}\in\mathbb{R}^{(N_{a}+N_{m})\times d}$. The spatial embeddings are generated by mapping the central positions of $N_{a}$ agents and $N_{m}$ lane vectors into a $d$-dimensional space, which are then added to $\bm{C}_{\rm sce}$. Then, this scene encoding is fed into another standard Transformer encoder to enhance agent-lane spatial interaction, producing the scene context feature $\bm{C}_{\rm ctx}\in\mathbb{R}^{(N_{a}+N_{m})\times d}$.
IV-C Deviation Feature Fusion
As discussed in Section III-C, features that describe deviations in the current vehicle state from the normal state can serve as indicators of long-tailed scenarios. Therefore, in the encoding process, we specifically model these deviation features from both the target vehicle and agent interaction perspectives, providing a comprehensive understanding of the deviation state of the surrounding traffic environment.
@@ IV-C Deviation Feature Fusion
We introduce deviation inputs including both target vehicle and agent interaction perspectives, as shown in Fig. 6. The individual deviation input $\bm{X}_{\rm dev}^{\rm ind}\in\mathbb{R}^{T_{o}\times 6}$ comprises several components. As illustrated in Fig. 7, let $\theta^{t}$ and $v^{t}$ represent the absolute heading angle and velocity of the target vehicle at time $t$, respectively. Additionally, let $\alpha^{t}$ denote the orientation of the vehicle’s displacement vector at time $t$. We employ a 6D descriptor to encapsulate the individual deviation feature $\bm{X}_{\rm dev}^{\rm ind}$, defined as $[\theta^{t}-\theta^{0},\theta^{t}-\alpha^{0},\theta^{t}-\alpha^{t},\sqrt{\frac{1}{t}\sum_{t}(\theta^{t}-\bar{\theta})^{2}},v^{t}-v^{0},\sqrt{\frac{1}{t}\sum_{t}(v^{t}-\bar{v})^{2}}]$. This descriptor comprises six components: the change in heading, the change in heading relative to the initial displacement orientation, the angle between heading and the current displacement orientation, the standard deviation of heading, the change in velocity, and the standard deviation of velocity. These metrics can be used to quantify directional changes and variability in motion and velocity, and are irrelevant to current positions or coordinates. The group deviation feature $\bm{X}_{\rm dev}^{\rm grp}$ can be formulated as a 2D vector: $[\frac{1}{N_{a}}\sum_{n=1}^{N_{a}}(v^{t}_{n}-v^{t}),\sqrt{\frac{1}{N_{a}}\sum_{n=1}^{N_{a}}(\theta^{t}_{n}-\bar{\theta}^{t})^{2}}]$, including the surrounding agents’ velocity relative to the target vehicle and group heading deviation.
@@
Both the individual and group deviation features are separately projected into high-dimensional spaces using MLPs. The resulting representations are then added to form a comprehensive deviation feature $\bm{X}_{\rm dev}\in\mathbb{R}^{T_{o}\times d}$. Subsequently, deviation fusion is performed by concatenating $\bm{X}_{\rm dev}$ and $\bm{X}_{\rm tgt}$, followed by an MLP-based fuser and a Transformer encoder layer to capture temporal dependencies. The final output deviation feature $\bm{C}_{\rm dev}$ is passed to the decoder along with $\bm{C}_{\rm ctx}$, providing insights into both scene contextual information and motion deviation status.
IV-D Dual Query-Based Trajectory Decoding
@@
With the advancement of query-based decoding techniques in trajectory prediction tasks, researchers have extensively explored the use of distinct queries to dynamically decode future trajectories [20, 21, 36]. In this paper, we propose to decode future trajectories using a mode query and dual future queries. The mode query $\bm{Q}_{m}\in\mathbb{R}^{K\times d}$ captures the diversity of different modes, supporting multimodal prediction. The dual future queries consist of a regular future query $\bm{Q}_{f}^{\rm regular}\in\mathbb{R}^{T_{f}\times d}$ and a tail future query $\bm{Q}_{f}^{\rm tail}\in\mathbb{R}^{T_{f}\times d}$. $\bm{Q}_{f}^{\rm regular}$ is designed to model the future dynamics of the target vehicle, while $\bm{Q}_{f}^{\rm tail}$ is intended to capture the deviation status of the current traffic scenario, thereby reflecting the long-tail characteristics of a trajectory sample. These queries are initialized in the decoder and then processed through multistream decoder blocks, allowing the model to simultaneously integrate both scene-level contextual information and vehicle deviation status.
IV-D1 Multistream Decoder Block
@@
We facilitate the simultaneous learning of contextual and deviation information through a multistream decoder block. The multistream decoder block is a crucial module in our model, designed to process and integrate multiple streams of information using attention mechanisms. In a single multistream decoder block, multiple distinct information streams are processed independently yet simultaneously, allowing for continuous processing within the same block.
@@
The computational flow of our multistream decoder block is formally described in Algorithm 1. The input of a multistream decoder block is one query and a stream of memory features, which can be heterogeneous in shape and type. In the first cross-modal fusion phase, each feature stream undergoes layer normalization before being processed through a multihead attention mechanism:
[…]
@@
$\bm{Q}_{\text{mlp}}=\sum_{k=1}^{K}G_{k}(\bm{Q})\cdot\text{Expert}_{k}(\bm{Q})$ (9)
where $G_{k}$ denotes learned gating weights of the $k$-th expert. Each expert is essentially an MLP, enabling conditional computation based on input charactersitics. The final fusion stage employs self-attention over the aggregated representation, followed by another MLP layer to project features into the target space.
[…]
$\text{LayerNorm}(\bm{Q}),$ (11)
@@
Input: Query Embedding $\bm{Q}$, Heterogeneous Feature Streams $S=\{\bm{S}_{i}\}_{i=1}^{n_{s}}$
@@
2:     $\bm{Q}_{\rm attn}$leftarrow MultiHeadAttention(      LayerNorm($\bm{Q}$), LayerNorm($\bm{S}_{i}$), LayerNorm($\bm{S}_{i}$))
3:     $\bm{Q}$leftarrow $\bm{Q}$ + Dropout($\bm{Q}_{\rm attn}$)
@@
7:  $\bm{Q}_{\rm attn}$leftarrow MultiHeadAttention(      LayerNorm($\bm{Q}$), LayerNorm($\bm{Q}$), LayerNorm($\bm{Q}$))
8:  $\bm{Q}$leftarrow $\bm{Q}$ + Dropout($\bm{Q}_{\rm attn}$)
9:  $\bm{Q}_{\rm mlp}$leftarrow MLP(LayerNorm($\bm{Q}$))
10:  $\hat{\bm{Q}}$leftarrow $\bm{Q}$ + Dropout($\bm{Q}_{\rm mlp}$)
@@
IV-D2 Dual Query-Based Trajectory Decoder
We propose decoding future trajectories in CDKFormer ultlizing the multistream decoder block introduced in Section IV-D1. The multistream decoder block is structured distinctly for each query type to accommodate both fine-grained scene understanding and long-tail trajectory prediction. For the mode query and the regular future query, the multistream decoder block operates through two distinct streams: the deviation feature $\bm{C}_{\rm dev}\in\mathbb{R}^{T\times d}$, which learns the motion and deviation patterns of the target vehicle, and the contextual encoding $\bm{C}_{\rm ctx}\in\mathbb{R}^{(N_{a}+N_{m})\times d}$, which sequentially captures scene semantics. This multistream fusion paradigm supports spatial-temporal knowledge learning. For the tail future query, the multistream decoder block processes only the deviation feature stream, aiming to primarily extract the motion and deviation status of the target vehicle. The dual future query design allows two queries to each carry out the task of learning regular and irregular traffic state dynamic, potentially enhancing long-tail trajectory oprediction. The calculation is shown below.

### v2

@@ 3.3 Long-Tail Characteristics
Utilizing Argoverse 2 motion forecasting dataset [wilson2023argoverse], we investigate the long-tail characteristics of the target vehicle trajectories from two perspectives: individual motion and group interaction. Our analysis focuses on comparing the top 10% of tail samples with the top 10% of head samples, based on the defined tail score. The following metrics are proposed to quantitatively measure the deviations in individual motion and group interactions between tailed and normal scenarios.
@@
These metrics, which represent the complexity of traffic scenarios surrounding the target vehicle, are independent of absolute positions. Consequently, they can be utilized to effectively assess the deviation of the current traffic state in a coordinate-agnostic way.
A descriptive table on these metrics and the results of significance comparisons from ANOVA tests are summarized in Table 1. The distributions of individual and group deviation measures are shown in Figure 3 and Figure 4, respectively. As a result, tail samples exhibit significantly higher mean changes in velocity and greater variability in both velocity and heading. For example, among the top 10% of tail samples, 26.74% involve clear turning maneuvers, while the remaining 73.26% proceed straight. In contrast, 88.28% of the top 10% head samples follow straight trajectories. Additionally, tail samples show more dynamic agent interactions with higher relative speed and greater group heading variability. These distinctions underscore the complexity and unpredictability inherent in long-tailed vehicle trajectories, highlighting the necessity of incorporating these features to fully understand and model long-tail traffic scenarios.
Table 1: Long-tail characteristics of trajectory samples in Argoverse 2 motion forecasting dataset.
@@
[…]
@@
4 Methodology
4.1 Prediction Framework Overview
The proposed CDKFormer framework is illustrated in Figure 5. CDKFormer is constructed in an encoder-decoder way. In the scene encoder, we first encode the HD map vectors and agent motion information and fuse them with a scene context fusion module. The deviation feature is learned in parallel with a deviation fusion module. Then, the learned context feature and deviation feature are jointly decoded in a query-based paradigm. A mode query and dual future queries are initiated to interactively and progressively extract the context and deviation feature through multistream decoder blocks. Then a scene query is obtained using a learnable gating mechanism. We additionally refine this scene query and generate the final multimodal trajectory predictions.
Figure 5: Overview of the proposed CDKFormer architecture. The model first encode the agent motion and scene contextual information with self-attention-based encoders. The deviation and motion features of the target vehicle are jointly fused in a deviation fusion module. The scene context and deviation information are subsequently decoded by a mode query and dual future queries, including a regular future query and a tail future query, within multistream decoder blocks. Then, a scene query is obtained by combining the mode query and weighted combined future query. This scene query is further refined and used for multimodal trajectories generation. $\times N$ denotes NN stacked layers. (R)Future and (T)Future denote regular future query and tail future query, respectively.
4.2 Scene Context Encoding
4.2.1 Agent Motion Encoding
The agent motion input is a 6-dimensional vector $\bm{X}\in\mathbb{R}^{N_{a}\times T_{o}\times d_{a}}$, where $d_{a}=6$ includes the 2D historical position, positional displacement vector, positional displacement magnitude, and absolute velocity. Fourier embedding [tancik2020fourier] is first applied to these motion inputs, followed by a multilayer perceptron (MLP) to project them into a high-dimensional space, producing the group motion encoding $\bm{X}_{a}\in\mathbb{R}^{N_{a}\times T_{o}\times d}$. Then, we apply a standard Transformer encoder on $\bm{X}_{a}$, which is composed of a multihead attention and an MLP, with layer normalization and residual connections added. This enables feature aggregation in the temporal dimension. The multihead attention is calculated as follows.
@@ 4.2.1 Agent Motion Encoding
where $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are weight matrix, $\bm{X}$ is the input, $d_{k}$ is the hidden state dimension. The encoder layer is repeated for $N$ times to obtain the motion feature $\bm{\tilde{X}}_{a}\in\mathbb{R}^{N_{a}\times T_{o}\times d}$, which represents the motion dynamic of both the target vehicle and its surrounding agents. Then we extract the target vehicle motion feature $\bm{X}_{\rm tgt}\in\mathbb{R}^{T_{o}\times d}$ from $\bm{\tilde{X}}_{a}$, which will be used for deviation feature fusion in Section 4.3. The final agent motion encoding is obtained by extracting the last timestep of $\bm{\tilde{X}}_{a}$, denoted as $\bm{C}_{a}\in\mathbb{R}^{N_{a}\times d}$.
4.2.2 Map Encoding
The input HD map feature $\bm{I}\in\mathbb{R}^{N_{m}\times l_{m}\times d_{m}}$ comprises the positions and displacement vectors of the centerlines surrounding the target vehicle, organized into $N_{m}$ polylines, each containing $l_{m}$ points with $d_{m}=4$. We encode these polylines with a PointNet-like encoder [qi2017pointnet], as also employed by [shi2022motion, cheng2023forecast].
@@ 4.2.2 Map Encoding
4.2.3 Scene Context Fusion
Scene context fusion is performed to fully integrate traffic agent motion and map information, enabling a semantic understanding of dynamic traffic environments. We first concatenate $\bm{C}_{a}\in\mathbb{R}^{N_{a}\times d}$ and $\bm{C}_{m}\in\mathbb{R}^{N_{m}\times d}$ to a scene encoding $\bm{C}_{\rm sce}\in\mathbb{R}^{(N_{a}+N_{m})\times d}$. Spatial embeddings are generated by mapping the central positions of $N_{a}$ agents and $N_{m}$ lane vectors into a $d$-dimensional space, which are then added to $\bm{C}_{\rm sce}$. Then, this scene encoding is fed into another standard Transformer encoder to enhance agent-lane spatial interaction, producing the scene context feature $\bm{C}_{\rm ctx}\in\mathbb{R}^{(N_{a}+N_{m})\times d}$.
4.3 Deviation Feature Fusion
As discussed in Section 3.3, features that describe deviations in the current vehicle state from the normal state can serve as indicators of long-tailed scenarios. Therefore, in the encoding process, we specifically model these deviation features from the perspective of the target vehicle and the agent interaction, providing a comprehensive understanding of the deviation state of the surrounding traffic environment.
@@ 4.3 Deviation Feature Fusion
We introduce deviation inputs including both target vehicle and agent interaction perspectives, as shown in Figure 6. The individual deviation input $\bm{X}_{\rm dev}^{\rm ind}\in\mathbb{R}^{T_{o}\times 6}$ comprises several components. As illustrated in Figure 7, let $\theta^{t}$ and $v^{t}$ represent the absolute heading angle and velocity of the target vehicle at time $t$, respectively. Furthermore, let $\alpha^{t}$ denote the orientation of the vehicle’s displacement vector at time $t$. We employ a 6D descriptor to encapsulate the individual deviation feature $\bm{X}_{\rm dev}^{\rm ind}$, defined as $\theta^{t}-\theta^{0},\theta^{t}-\alpha^{0},\theta^{t}-\alpha^{t},\sqrt{\frac{1}{t}\sum_{t}(\theta^{t}-\bar{\theta})^{2}},v^{t}-v^{0},\sqrt{\frac{1}{t}\sum_{t}(v^{t}-\bar{v})^{2}}]$, where $\bar{\theta}$ and $\bar{v}$ are the mean heading and mean velocity of the target vehicle over the observation window, respectively. This descriptor comprises six components: the change in heading, the change in heading relative to the initial displacement orientation, the angle between heading and the current displacement orientation, the standard deviation of heading, the change in velocity, and the standard deviation of velocity. These metrics can be used to quantify directional changes and variability in motion and velocity, and are irrelevant to current positions or coordinates. The group deviation feature $\bm{X}_{\rm dev}^{\rm grp}$ can be formulated as a 2D vector: $\frac{1}{N_{a}}\sum_{n=1}^{N_{a}}(v^{t}_{n}-v^{t}),\sqrt{\frac{1}{N_{a}}\sum_{n=1}^{N_{a}}(\theta^{t}_{n}-\bar{\theta}^{t})^{2}}]$, including the surrounding agents’ velocity relative to the target vehicle and group heading deviation, where $\theta_{n}^{t}$ is the heading of the n-th surrounding agent at time t, and $\bar{\theta}^{t}$ is the mean heading of all surrounding agents at time t.
@@
Individual and group deviation features are separately projected into high-dimensional spaces using MLPs. The resulting representations are then added to form a comprehensive deviation feature $\bm{X}_{\rm dev}\in\mathbb{R}^{T_{o}\times d}$. Subsequently, deviation fusion is performed by concatenating $\bm{X}_{\rm dev}$ and $\bm{X}_{\rm tgt}$, followed by an MLP-based fuser and a Transformer encoder layer to capture temporal dependencies. The final output deviation feature $\bm{C}_{\rm dev}$ is passed to the decoder along with $\bm{C}_{\rm ctx}$, providing insights into both scene contextual information and motion deviation status.
4.4 Dual Query-Based Trajectory Decoding
@@
With the advancement of query-based decoding techniques in trajectory prediction tasks, researchers have extensively explored the use of distinct queries to dynamically decode future trajectories [shi2022motion, zhou2023query, zhang2024decoupling]. In this paper, we propose decoding future trajectories using a mode query and dual future queries. The mode query $\bm{Q}_{m}\in\mathbb{R}^{K\times d}$ captures the diversity of different modes, which supports multimodal prediction. Dual future queries consist of a regular future query $\bm{Q}_{f}^{\rm regular}\in\mathbb{R}^{T_{f}\times d}$ and a tail future query $\bm{Q}_{f}^{\rm tail}\in\mathbb{R}^{T_{f}\times d}$. $\bm{Q}_{f}^{\rm regular}$ is designed to model the future dynamics of the target vehicle, while $\bm{Q}_{f}^{\rm tail}$ is intended to capture the deviation status of the current traffic scenario, thus reflecting the long tail characteristics of a trajectory sample. These queries are initialized as learnable embedding vectors in the decoder and then processed through multistream decoder blocks, allowing the model to simultaneously integrate both scene-level contextual information and vehicle deviation status.
4.4.1 MultiStream Decoder Block
@@
We facilitate simultaneous learning of contextual and deviation information through a multistream decoder block. The multistream decoder block is a crucial module in our model, designed to process and integrate multiple streams of information using attention mechanisms. As shown in Figure 8, in a single multistream decoder block, multiple distinct information streams are processed independently yet simultaneously, allowing continuous processing within the same block.
Figure 8: MultiStream Decoder Block. (a) The structure of the MultiStream Decoder Block. (b) The structure of the MoE layer in cross-modal fusion phase.
@@
The computational flow of our multistream decoder block is formally described in Algorithm 1. The input of a multistream decoder block is one query and a stream of $K_{s}$ memory features, which can be heterogeneous in shape and type. In the first cross-modal fusion phase, each feature stream undergoes layer normalization before being processed through a multihead attention mechanism:
$\bm{Q}_{\text{attn}}=\text{MultiHeadAttention}($ $\text{LayerNorm}(\bm{Q}),$ (7)
@@
$\bm{Q}_{\text{mlp}}=\sum_{k=1}^{K_{e}}G_{k}(\bm{Q})\cdot\text{Expert}_{k}(\bm{Q})$ (10)
where the expert networks and the gating function are defined as:
$\text{Expert}_{k}(\bm{Q})=\text{ReLU}(\bm{Q}\bm{W}_{k1}+\bm{b}_{k1})\bm{W}_{k2}+\bm{b}_{k2}$
$G(\bm{Q})=\text{Softmax}(\bm{Q}\bm{W}_{g}+\bm{b}_{g})$
where $\bm{G}(\bm{Q})$ is the gating weights from the softmax output, and $G_{k}(\bm{Q})$ is its k-th component. Each expert is a two-layer MLP that processes the query, and $K_{e}$ is the total number of experts. $\bm{W}_{k1},\bm{b}_{k1},\bm{W}_{k2},\bm{b}_{k2},\bm{W}_{g},\bm{b}_{g}$ are learnable weights.
The final fusion stage employs self-attention over the aggregated representation, followed by another MLP layer to project features into the target space.
$\hat{\bm{Q}}=\text{MultiHeadAttention}($ $\text{LayerNorm}(\bm{Q}),$
$\text{LayerNorm}(\bm{Q}),$
@@
Input: Query Embedding $\bm{Q}$, Heterogeneous Feature Streams $S=\{\bm{S}_{i}\}_{i=1}^{K_{s}}$
@@
2:  $\bm{Q}_{\rm attn}$ $\leftarrow$ MultiHeadAttention(   LayerNorm($\bm{Q}$), LayerNorm($\bm{S}_{i}$), LayerNorm($\bm{S}_{i}$))
3:  $\bm{Q}$ $\leftarrow$ $\bm{Q}$ + Dropout($\bm{Q}_{\rm attn}$)
@@
[…]
@@
4.4.2 Dual Query-Based Trajectory Decoder
We propose decoding future trajectories in CDKFormer ultlizing the multistream decoder block introduced in Section 4.4.1. The multistream decoder block is structured distinctly for each query type to accommodate both fine-grained scene understanding and long-tail trajectory prediction. For the mode query and the regular future query, the multistream decoder block operates through two distinct streams: the deviation feature $\bm{C}_{\rm dev}\in\mathbb{R}^{T\times d}$, which learns the motion and deviation patterns of the target vehicle, and the contextual encoding $\bm{C}_{\rm ctx}\in\mathbb{R}^{(N_{a}+N_{m})\times d}$, which sequentially captures scene semantics. This multistream fusion paradigm supports spatial-temporal knowledge learning. For the tail future query, the multistream decoder block processes only the deviation feature stream, aiming to primarily extract the motion and deviation status of the target vehicle. The dual future query design allows two queries to each perform the task of learning regular and irregular traffic state dynamic, potentially enhancing long-tail trajectory prediction. The calculation is shown below.